OpenStudy (anonymous):

Instruction:Solve each equation. Then check your solution!! 1/4(5-2h)=g/6

OpenStudy (anonymous):

you have two variables, \(g\) and \(h\) . which one are you supposed to solve for?

OpenStudy (anonymous):

sorry the question: 1/4(5-2h)=h/2

OpenStudy (anonymous):

make your life easier by multiplying both sides by 4 first to clear the fractions \[\frac{1}{4}(5-2h)=\frac{h}{2}\] \[4\times \frac{1}{4}(5-2h)=\frac{h}{2}\times4\] \[5-2h=2h\] as a start

OpenStudy (anonymous):

then add \(2h\) to both sides, and finally divide both sides by \(4\)

OpenStudy (anonymous):

your question had \(\frac{g}{6}\) on the right and not \(\frac{h}{2}\) if it is \(\frac{h}{6}\) then the answer i wrote above is incorrect. you would need to multiply both sides by \(12\) to clear the fractions

OpenStudy (anonymous):

can u like work the problem and show the steps!!

OpenStudy (anonymous):

are u there??

OpenStudy (anonymous):

i did work the problem and did show the steps above

OpenStudy (anonymous):

wat do u mean!! I should 2h to both sides then divide each side by 4!! I don't get that part!!